# Structural barriers to complete homogenization and wormholing in dissolving porous and fractured rocks

### v1

@@ 3.3 Discrete Fracture Network
Figure 7: Evolution of the discrete fracture network and the flow focusing profile in three dissolution regimes: (a) uniform ($\textrm{Da}_{\textrm{eff}}=0.0002$, $\textrm{G}=5$), (b) channeling ($\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$), and (c) wormholing ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$). The network is shown as a graph after a principal-component-analysis projection; edge width scales with flow, but only edges carrying more than 1% of the maximum flow rate in the system are plotted.
@@
Figure 8: Evolution of the flow focusing profile for the discrete fracture network across the ($\textrm{Da}_{\textrm{eff}}$, G) parameter space. (a) Mean profile value at successive time snapshots (solid lines). (b) Same data, with ensemble variability: the shaded band around each curve denotes the mean $\pm$ 1 standard deviation across all realizations.
@@
Figure 7 presents the evolution of the dissolving DFNs and Figure8 depicts the corresponding changes in the flow focusing profile across the parameter space. We perform the simulations for $\textrm{Da}_{\textrm{eff}}\in\{0.0002,0.002,0.02\}$ and $\textrm{G}\in\{0.1,1,5\}$, and capture data at $T\in\{0,0.1,0.2,0.5,1\}$. The initial values of the flow focusing index are very high ($f_{50\%}\approx 0.8$), indicating that the network is strongly channelized even before any reaction occurs. Because the discrete fracture network has a much larger hydraulic cross-section than the pore lattices, the effective Damköhler numbers required to traverse the three dissolution regimes are shifted roughly one order of magnitude lower. After accounting for this shift, the expected progression—uniform dissolution, channeling, and finally wormholing—emerges unchanged.
@@
In the uniform regime the flow focusing profile shows a modest decrease, maximally to around $f_{50\%}\approx 0.7$, but the values remain high, showing that uniform aperture growth cannot eliminate heterogeneity rooted in fracture connectivity and pathway lengths, making complete homogenization unattainable in a DFN.
@@
In the channeling regime the flow focusing profile rises only slightly with time, as expected for a network that is already strongly channelized. Its shape, however, evolves much like in porous media: the increase is nearly uniform across the entire domain. In addition, for $\textrm{Da}_{\textrm{eff}}=0.002$ and $G=0.1$, the DFN transitions to the wormholing regime, mirroring the behavior of the disordered pore network and further underscoring the influence of additional heterogeneity.
@@
In the wormholing regime the flow focusing profile develops a transition zone, much like in porous media. A notable difference, however, is that a higher G slows wormhole advance—the opposite of the behavior seen in pore networks. At the same time, the flow focusing index no longer approaches its maximum when G is large. This indicates that diffusion limitation becomes significant, restricting dissolution in the main channel. As a result, competition among flow paths weakens; smaller channels continue to carry a portion of the flux, thereby reducing the growth rate of the dominant pathway.
@@
The variability in the flow focusing index (Figure 8b) mirrors that of the disordered pore network, rising at first and then declining as dissolution proceeds. In the DFNs the wormholes ultimately reach the outlet, and, once breakthrough occurs, the profile exhibits the expected reduction in variability.
@@ 3.4 Homogenization by Dissolution
Figure 9: Velocity distributions for three network types: (a) regular pore network, (b) disordered pore network, and (c) discrete fracture network. Each panel contains three histograms: initial distribution, after uniform dissolution, and the distribution for the same ensemble, but with uniform diameters or apertures. Velocities are normalized by the ensemble-average velocity in each case.
The flow focusing data show that perfect homogenization is unattainable in either disordered porous media or DFNs. To examine why, we compare the velocity distributions for all three network types. Specifically, we sample the initial velocity distribution, the distribution after uniform dissolution (final snapshots for $\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$ in Figures 4 and 6, and $\textrm{Da}_{\textrm{eff}}=0.0002$, $\textrm{G}=5$ in Figure 8), and the distribution for an ideal network with perfectly uniform diameters and apertures.
Figure 9 presents the comparison of these distributions. For the regular pore network, the initial histogram is broad because of diameter variability; after dissolution it narrows sharply and collapses to a single value in the uniform diameter limit. This behavior is consistent with other studies that use the regular pore network [63, 15].
In the disordered network the initial velocity histogram is broad, with a high-velocity tail generated by the largest pores. Uniform dissolution trims this tail by reducing the contrast between large and small pores. At the same time, very low velocities become more common: as heterogeneity diminishes, flow aligns with the main pressure gradient and many pores oriented perpendicular to it carry negligible flux. Interestingly, the histogram after dissolution almost coincides with that of the uniform‐diameter lattice, confirming that the remaining heterogeneity—and thus the residual flow focusing index—is rooted in network structure, not in diameter variance.
@@ 4 Conclusions
Dissolution reshapes both structure and flow in porous and fractured media – systems whose contrasting architectures and transport behaviors influence many geologic and engineering processes. To quantify these effects, we performed network-based simulations on three models: a regular pore network, a disordered pore network, and a discrete fracture network, spanning a wide range of flow and reaction conditions. The results are quantified using the flow focusing profile. Every dissolution regime presents distinct characteristics in each system. In the uniform regime, all networks are homogenized and flow focusing decreases across the entire domain. However, while the flow in the regular medium becomes completely uniform, the disordered medium retains a certain level of focusing, and the discrete fracture network remains highly heterogeneous, with only minimal redirection of flow to the parts of the network beyond the main pathways.
@@
We focused our analyses on the interplay between heterogeneity and uniform dissolution. The three systems exhibited different types and degrees of heterogeneity, and each responded to dissolution in its own way. In the regular network, heterogeneity arose solely from the pore-diameter distribution and was readily eliminated by uniform dissolution. In the disordered network and the DFN, additional structural heterogeneity persisted: the lengths of individual pathways and the connectivity between parts of the network remained unchanged. Uniform dissolution therefore removed only the heterogeneity associated with diameter or aperture variations, demonstrating that the inherent structure exerted strong control over dissolution dynamics.
These findings also carried important implications for continuum-scale modeling. In our simulations dissolution could eradicate heterogeneity in conduit diameters (or apertures) but could not remove heterogeneity rooted in pathway lengths and network connectivity; as a result, flow never became fully uniform in the disordered lattice or in the DFN. Darcy-scale finite-difference models, however, represent the medium only through spatially varying porosity or permeability fields. Because those fields evolve toward uniform values when diameters enlarge uniformly, such models would predict nearly complete flow homogenization under uniform dissolution—an outcome that, according to our network results, is unattainable in natural media where connectivity and length distributions persist. Capturing the true limits of homogenization will therefore require numerical approaches that retain at least some information on network topology, for example through higher-order upscaling techniques or hybrid continuum–network formulations.
@@ 5 Data Availability
All simulation data files are available in [71]. The software for dissolution and flow focusing profile calculation is provided in [72].
@@ Appendix A DFN Generation
We generate three-dimensional discrete fracture networks (DFN) using the dfnWorks [37] software suite. The cornerstone of the DFN methodology, is that fractures tend to have a radius (length) that is much larger than their aperture (height). This disparity is used to justify a co-dimension one representation of each fracture. For example, each fracture is a one-dimensional line in a two-dimensional simulation or a two-dimensional plane in a three-dimensional simulation. Thus, in our 3D DFN model, each fracture is represented as a rectangle using the methods described in [34]. These planes intersect with one another to form a network, through which flow and transport is simulated. The size, shape, orientation, and other hydrological properties are sampled from distributions whose parameters are determined from a site characterization, cf. [78] for a comprehensive discussion of DFN modeling approaches. Thus, dfnWorks stochastically generates three-dimensional DFNs with desired characteristics such as length distributions and fracture intensities. Details of dfnWorks in terms of algorithms and various applications can be found in [37].
We consider a semi-generic DFN loosely based on the fractured carbonate-hosted Pietrasecca Fault in the central Apennines, Italy, cf. [66] for detailed information about the fractured media. The domain is 25 m $\times$ 10 m $\times$ 10 m. We adopt three families from the survey and modify a few of their properties for our semi-generic study. Generation parameters of the networks are provided in Table 1. We measure network surface area in terms of the fracture intensity [m-1], which is the sum of surface area of each fracture ($S_{f}$) divided by the volume of the domain,
@@
We modify the minimum and maximum fracture radius from the data provided in [66]. This modification requires a change in the values of fracture intensity $P_{32}$. Fracture family orientations are sampled from a three dimensional von Mises Fisher distribution,
@@
In (9), $\boldsymbol{\mu}$} is the mean direction vector of the fracture family, $T$ denotes transpose, and $\kappa\geq 0$ is the concentration parameter that determines the degree of clustering around the mean direction. Values of $\kappa$ close to zero lead to a uniform distribution of points on the sphere while larger values create points with a small deviation from mean direction. To obtain uniformly random orientations, we set $\kappa=0.1$ and mean normal vector of $(0,0,1)$. The distribution is sampled using the method detailed in [79].
Discrete Fracture Network Families
@@ A.1 Graph Representation
Once the networks are generated, we derive a graph / pipe-network representation using the method presented in [35]. At the core of the DFN methodology is the conceptual model of a set of fractures, which are discrete entities intersecting with one another to form a network. The following maps a DFN $\mathcal{F}$ made up of a set of $n$ fractures $\{f_{i}\}$ to a graph $\mathcal{G}$ composed of a set of vertices $V$ and edges $E$. If two fractures $f_{i}$ and $f_{j}$ intersect, $f_{i}\cap f_{j}\neq\emptyset$, then there is a vertex $v\in V$,
$\psi:f_{i}\cap f_{j}\rightarrow v.$ (10)

### v2

@@ 3.3 Discrete Fracture Network
Figure 7: Evolution of the discrete fracture network and the flow focusing profile in three dissolution regimes: (a) uniform ($\textrm{Da}_{\textrm{eff}}=0.0002$, $\textrm{G}=5$), (b) channeling ($\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$), and (c) wormholing ($\textrm{Da}_{\textrm{eff}}=0.02$, $\textrm{G}=5$). The network is shown as a graph after a principal-component-analysis projection; only edges carrying more than 1% of the maximum flow rate in the system are plotted. Edge width is proportional to the volumetric flow rate, with the proportionality constant rescaled to maximize visibility but kept fixed within each regime. The plots show the initial flow focusing profile (black line) and profiles at times $T=0.1,0.2,0.5,$ and $1.0$ (colored lines). The frame color denotes the dissolution regime: green for uniform, red for channeling, and blue for wormholing.
@@
Figure 8: Evolution of the flow focusing profile for the discrete fracture network across the ($\textrm{Da}_{\textrm{eff}}$, G) parameter space. (a) Mean profile value at successive time snapshots (solid lines). (b) Same data, with ensemble variability: the shaded band around each curve denotes the mean $\pm$ 1 standard deviation across all realizations. The plots show the initial flow focusing profile (black line) and profiles at times $T=0.1,0.2,0.5,$ and $1.0$ (colored lines). The frame color denotes the dissolution regime: green for uniform, red for channeling, and blue for wormholing.
@@
Fig. 7 presents the evolution of the dissolving DFNs and Fig. 8 depicts the corresponding changes in the flow focusing profile across the parameter space. We perform the simulations for $\textrm{Da}_{\textrm{eff}}\in\{0.0002,0.002,0.02\}$ and $\textrm{G}\in\{0.1,1,5\}$, and capture data at $T\in\{0.0,0.1,0.2,0.5,1.0\}$. The initial values of the flow focusing index are very high ($f_{50\%}\approx 0.8$ in Fig. 8), indicating that the network is strongly channelized even before any reaction occurs. Because the discrete fracture network has a much larger hydraulic cross-section than the pore lattices, the effective Damköhler numbers required to traverse the three dissolution regimes are shifted roughly one order of magnitude lower. After accounting for this shift, the expected progression—uniform dissolution, channeling, and finally wormholing—emerges unchanged.
@@
As visible in the top rows of Fig. 8a,b, in the uniform regime the flow focusing profile shows a modest decrease, maximally to around $f_{50\%}\approx 0.7$ in the top-right panels of Fig. 8a,b, but the values remain high, showing that uniform aperture growth cannot eliminate heterogeneity rooted in fracture connectivity and path lengths, making complete homogenization unattainable in a DFN.
@@
In the channeling regime (middle and middle-right panels of Fig. 8a,b), the flow focusing profile rises only slightly with time, as expected for a network that is already strongly channelized. Its shape, however, evolves much like in porous media: the increase is nearly uniform across the entire domain. In addition, for $\textrm{Da}_{\textrm{eff}}=0.002$ and $\textrm{G}=0.1$, the DFN transitions to the wormholing regime, mirroring the behavior of the disordered pore network and further underscoring the influence of additional heterogeneity.
@@
In the wormholing regime (middle-left panels and bottom rows of Fig. 8a,b), the flow focusing profile develops a transition zone, much like in porous media. A notable difference, however, is that a higher G slows wormhole advance—the opposite of the behavior seen in pore networks. At the same time, the flow focusing index no longer approaches its maximum when G is large. This indicates that diffusion limitation becomes significant, restricting dissolution in the main channel. As a result, competition among flow paths weakens; smaller channels continue to carry a portion of the flux, thereby reducing the growth rate of the dominant path.
@@
The variability in the flow focusing index (Fig. 8b) mirrors that of the disordered pore network, rising at first and then declining as dissolution proceeds. In the DFNs the wormholes ultimately reach the outlet, and, once breakthrough occurs, the profile exhibits the expected reduction in variability.
@@ 3.4 Homogenization by Dissolution
Figure 9: Velocity distributions for three network types: (a) regular pore network, (b) disordered pore network, and (c) discrete fracture network. Each panel shows three histograms: the initial distribution (black), the distribution after uniform dissolution (blue), and the distribution for the same ensemble but with uniform diameters or apertures (orange). Velocities are normalized by the ensemble-average velocity in each case.
The flow focusing data show that perfect homogenization is unattainable in either disordered porous media or DFNs. To examine why, we compare the velocity distributions for all three network types. Specifically, we sample the initial velocity distribution, the distribution after uniform dissolution (final snapshots for $\textrm{Da}_{\textrm{eff}}=0.002$, $\textrm{G}=5$ in Figs. 4 and 6, and $\textrm{Da}_{\textrm{eff}}=0.0002$, $\textrm{G}=5$ in Fig. 8), and the distribution for an ideal network with perfectly uniform diameters and apertures.
Fig. 9 presents the comparison of these distributions. For the regular pore network, the initial histogram is broad because of diameter variability; after dissolution it narrows sharply and collapses to a single value in the uniform diameter limit. This behavior is consistent with other studies that use the regular pore network. Roded et al. [2020], Deng et al. [2025]
In the disordered network the initial velocity histogram is broad, with a high-velocity tail generated by the largest pores. Uniform dissolution trims this tail by reducing the contrast between large and small pores. At the same time, very low velocities become more common: as heterogeneity diminishes, flow aligns with the main pressure gradient and many pores oriented perpendicular to it carry negligible flux. Interestingly, the histogram after dissolution almost coincides with that of the uniform‐diameter lattice, confirming that the remaining heterogeneity—and thus the residual flow focusing—is rooted in network structure, not in diameter variance.
@@ 4 Conclusions
Dissolution reshapes both structure and flow in porous and fractured media—systems whose contrasting architectures and transport behaviors influence many geologic and engineering processes. To quantify these effects, we perform network-based simulations on three models: a regular pore network, a disordered pore network, and a discrete fracture network, spanning a wide range of flow and reaction conditions. The results are quantified using the flow focusing profile. Every dissolution regime presents distinct characteristics in each system. In the uniform regime, all networks are homogenized and flow focusing decreases across the entire domain. However, while the flow in the regular medium becomes completely uniform, the disordered medium retains a certain level of focusing, and the discrete fracture network remains highly heterogeneous, with only minimal redirection of flow into parts of the network beyond the main paths.
@@
We focus our analyses on the interplay between heterogeneity and uniform dissolution. The three systems exhibit different types and degrees of heterogeneity, and each responds to dissolution in its own way. In the regular network, heterogeneity arises solely from the pore-diameter distribution and is readily eliminated by uniform dissolution. In the disordered network and the DFN, additional structural heterogeneity persists: the lengths of individual paths and the connectivity between parts of the network remain unchanged. Uniform dissolution therefore removes only the heterogeneity associated with diameter or aperture variations, demonstrating that the inherent structure exerts strong control over dissolution dynamics.
These findings carry important implications for continuum-scale modeling. In our simulations, dissolution can eradicate heterogeneity in conduit diameters (or apertures) but cannot remove heterogeneity rooted in path lengths and network connectivity; as a result, flow never becomes fully uniform in the disordered lattice or in the DFN. Darcy-scale finite-difference models, however, represent the medium only through spatially varying porosity or permeability fields. Because those fields evolve toward uniform values when diameters enlarge uniformly, such models would predict nearly complete flow homogenization under uniform dissolution—an outcome that, according to our network results, is unattainable in natural media where connectivity and length distributions persist. Capturing the true limits of homogenization will therefore require numerical approaches that retain at least some information on network topology, for example through higher-order upscaling techniques or hybrid continuum–network formulations.
Our work extends the application of the flow focusing profile from disordered pore networks Szawełło et al. [2024] to a broader class of systems, including regular networks and discrete fracture networks, and shows that conduit-, segment-, and network-scale heterogeneity each leave a distinct imprint on dissolution regimes. The metric provides a unified way to compare porous and fractured media across parameter space and to identify when channeling, wormholing, or uniform dissolution are limited by quenched structural disorder. The strongly varying levels of this disorder in porous media and fracture networks indicate that care is needed when extrapolating laboratory experiments—often based on core dissolution in porous media—to field-scale fracture networks. The emergent properties of the dissolving medium, and its transport behavior, may differ substantially from those inferred from such experiments. This implies that predictions of stimulation efficiency, injectivity evolution, or reactive-front propagation in applications such as acidizing, geothermal operations, and geological carbon storage must account for the structural heterogeneity encoded in path lengths and connectivity. Designs or models that assume uniform dissolution can fully homogenize flow are likely to underestimate the persistence of preferential paths, and therefore the degree of channelized transport, even in regimes that appear uniform at the pore scale.
@@ 5 Data Availability
All simulation data files are available in [dataset] Szawełło et al. Szawełło et al. [2025] The software for dissolution and flow focusing profile calculation is provided in Szawełło. Szawełło [2025]
@@ Appendix A DFN Generation
We generate three-dimensional discrete fracture networks (DFN) using the dfnWorks Hyman et al. [2015] software suite. The cornerstone of the DFN methodology is that fractures tend to have a radius (length) that is much larger than their aperture (height). This disparity is used to justify a co-dimension one representation of each fracture. For example, each fracture is a one-dimensional line in a two-dimensional simulation or a two-dimensional plane in a three-dimensional simulation. Thus, in our 3D DFN model, each fracture is represented as a rectangle using the methods described in Hyman et al. Hyman et al. [2014] These planes intersect with one another to form a network, through which flow and transport is simulated. The size, shape, orientation, and other hydrological properties are sampled from distributions whose parameters are determined from a site characterization, cf. Viswanathan et al. Viswanathan et al. [2022] for a comprehensive discussion of DFN modeling approaches. Thus, dfnWorks stochastically generates three-dimensional DFNs with desired characteristics such as length distributions and fracture intensities. Details of dfnWorks in terms of algorithms and various applications can be found in Hyman et al. Hyman et al. [2015]
We consider a semi-generic DFN loosely based on the fractured carbonate-hosted Pietrasecca Fault in the central Apennines, Italy, cf. Smeraglia et al. Smeraglia et al. [2021] for detailed information about the fractured media. The domain is 25 m $\times$ 10 m $\times$ 10 m. We adopt three families from the survey and modify a few of their properties for our semi-generic study. Generation parameters of the networks are provided in Table 1. We measure network surface area in terms of the fracture intensity [m-1], which is the sum of surface area of each fracture ($S_{f}$) divided by the volume of the domain,
@@
We modify the minimum and maximum fracture radius from the data provided in Smeraglia et al. Smeraglia et al. [2021] This modification requires a change in the values of fracture intensity $P_{32}$. Fracture family orientations are sampled from a three dimensional von Mises Fisher distribution,
@@
In (14), $\boldsymbol{\mu}$} is the mean direction vector of the fracture family, $T$ denotes transpose, and $\kappa\geq 0$ is the concentration parameter that determines the degree of clustering around the mean direction. Values of $\kappa$ close to zero lead to a uniform distribution of points on the sphere while larger values create points with a small deviation from mean direction. $\bf x$} is a random 3-dimensional unit vector within entries sampled from a uniform distribution on $[0,1]$. Trend and Plunge are converted to a 3D normal vector using the standard convention where an upward-pointing normal (negated from the standard downward pole) is given by
${\boldsymbol{\mu}}[0]$ $=-\sin(\textrm{Trend})\cos(\textrm{Plunge}),$
${\boldsymbol{\mu}}[1]$ $=-\cos(\textrm{Trend})\cos(\textrm{Plunge}),$ (15)
${\boldsymbol{\mu}}[2]$ $=\sin(\textrm{Plunge}).$
To obtain uniformly random orientations, we set $\kappa=0.1$ and mean normal vector of $(0,0,1)$. The distribution is sampled using the method detailed in Wood. Wood [1994]
DFN Families
@@ A.1 Graph Representation
Once the networks are generated, we derive a graph/pipe-network representation using the method presented in Hyman et al. Hyman et al. [2018] At the core of the DFN methodology is the conceptual model of a set of fractures, which are discrete entities intersecting with one another to form a network. The following maps a DFN $\mathcal{F}$ made up of a set of $n$ fractures $\{f_{i}\}$ to a graph $\mathcal{G}$ composed of a set of vertices $V$ and edges $E$. If two fractures $f_{i}$ and $f_{j}$ intersect, $f_{i}\cap f_{j}\neq\emptyset$, then there is a vertex $v\in V$,
$\psi:f_{i}\cap f_{j}\rightarrow v$ (16)